An object accelerates uniformly in a straight line at a rate of 0.250 m/s2 from 1.50 m/s to 4.00 m/s. How far did the object travel during this acceleration in units of meter? (don't write the unit


1
Expert's answer
2021-10-07T08:49:33-0400

The distance can be found with a very useful formula:


"d=\frac{v_f^2-v_i^2}{2a}=\frac{4^2-1.5^2}{2·0.25}=27.5\text{ m}."